OpenStudy (anonymous):

If f(x) = 1/3x − 2 and g(x) = −2x − 3, find (f • g)(6)

OpenStudy (anonymous):

(f.g)(6)=f(6).g(6)

OpenStudy (anonymous):

i got f=96 and g=-90 is this right?

OpenStudy (anonymous):

\[f\cdot g\] or \[f\circ g\]

OpenStudy (anonymous):

sorry im confused

OpenStudy (anonymous):

g(6) = -2*6 - 3 = -12 -3 = -15

OpenStudy (anonymous):

i dont know how you came to that answer. i must not know how to work the problem. or we are doing two different ways because I got the same answer.

OpenStudy (anonymous):

I just plugged 6 into the equation for x.

OpenStudy (anonymous):

okay thank you.

OpenStudy (anonymous):

welcome